### v1

Abstract
While Large Language Models (LLMs) have demonstrated remarkable potential in time series forecasting, their practical deployment remains constrained by excessive computational demands and memory footprints. Existing LLM-based approaches typically suffer from three critical limitations: (1) Inefficient parameter utilization in handling numerical time series patterns; (2) Modality misalignment between continuous temporal signals and discrete text embeddings; and (3) Inflexibility for real-time expert knowledge integration. We present Small but Mighty Enhancing Time Series (SMETimes), the first systematic investigation of sub-3B parameter Small Language Models (SLMs) for efficient and accurate time series forecasting. Our approach centers on three key innovations: (1) A statistically-enhanced prompting mechanism that bridges numerical time series with textual semantics through descriptive statistical features; (2) A adaptive fusion embedding architecture that aligns temporal patterns with language model token spaces through learnable parameters; And (3) a dynamic mixture-of-experts framework enabled by SLMs’ computational efficiency, adaptively combining base predictions with domain-specific models. Extensive evaluations across seven benchmark datasets (ETTh1/2, ETTm1/2, Weather, Solar, ECL) demonstrate that our 3B-parameter SLM achieves state-of-the-art performance on five primary datasets while maintaining 3.8$\times$ faster training and 5.2$\times$ lower memory consumption compared to 7B-parameter LLM baselines. Notably, the proposed model exhibits better learning capabilities, achieving 12.3% lower MSE than conventional LLM. Ablation studies validate that our statistical prompting and cross-modal fusion modules respectively contribute 15.7% and 18.2% error reduction in long-horizon forecasting tasks. By redefining the efficiency-accuracy trade-off landscape, this work establishes SLMs as viable alternatives to resource-intensive LLMs for practical time series forecasting. Code and models are available at https://github.com/xiyan1234567/SMETimes.
keywords:
Small Language Models, Statistically-enhanced Prompting, Adaptive Fusion Embedding
1 Introduction
Time series forecasting stands as a cornerstone of modern decision-making systems, with critical applications spanning energy management (petrucci2022modelling, ), financial markets (leonard2001promotional, ), climate modeling (schneider1974climate, ), and intelligent transportation (li2015demand, ). Traditional approaches often rely on domain-specific statistical models (box1994time, ) or deep neural networks (liu2023itransformer, ), which require substantial computational resources and extensive domain expertise. While large language models (LLMs) have recently demonstrated remarkable capabilities in time series forecasting (jin2023time, ; zhou2023ptse, ), their practical deployment faces significant challenges due to prohibitive computational costs and memory footprints (touvron2023llama, ).
Figure 1: Performance-efficiency Trade-off Comparison on ETTh1 (zhou2021informer, ) Dataset. Our SLM variants (blue) achieve competitive MSE with significantly lower training time and memory footprint compared to conventional LLM-based approaches. Bubble size represents relative memory consumption.
The recent proliferation of LLM-based forecasting methods (woo2024unified, ; gruver2023large, ) has revealed three fundamental limitations: (1) Massive parameter counts (typically $>$7B) lead to inefficient training/inference; (2) Inadequate alignment between numerical time series patterns and textual embeddings; And (3) limited flexibility for integrating domain-specific expert knowledge. As shown in Figure 1, conventional LLM approaches like Time-LLM (jin2023time, ) (22.33G) and AutoTimes liu2025autotimes (23.12G) incur substantial resource costs despite comparable MSE performance to our 1B-parameter SLM (4.33G). This efficiency gap becomes particularly critical in real-world deployment scenarios with hardware constraints.
To alleviate the computational inefficiency of large language models (LLMs) in time series forecasting, we propose a feature fusion strategy through Small Language Models (SLMs) that strategically reduce model scale while incorporating targeted architectural innovations to achieve better efficiency. Our key insight lies in three synergistic components: (1) Statistically-enhanced prompting that bridges numerical and textual domains, (2) A adaptive fusion mechanism for time series embeddings, And (3) dynamic mixture-of-experts (MoE) integration enabled by SLMs’ lightweight architecture. Extensive experiments across seven benchmark datasets (ETTh1/2 (zhou2021informer, ), ETTm1/2 (zhou2021informer, ), Weather (wu2021autoformer, ), Solar (lai2018modeling, ), ECL (wu2021autoformer, )) demonstrate that our 1B-parameter SLM achieves state-of-the-art results on five primary datasets while maintaining competitive performance on the remaining three.
Our work makes four fundamental contributions to the field of efficient time series forecasting:
  • To the best of our knowledge, this is the first work to develop and evaluate a framework that applies small language models (SLMs) to time series forecasting tasks. Through targeted architectural modifications, we demonstrate that compact models achieve performance parity with 7B-parameter large language models while attaining 3.2$\times$ accelerated training convergence and 5.1$\times$ reduced memory footprint, substantially enhancing deployment feasibility in resource-constrained environments.
  • We introduce a new prompting methodology integrating temporal statistics with domain-specific contextual metadata. Empirical validation reveals this statistically-enhanced approach reduces mean squared error by 12.7% compared to conventional textual prompting baselines through systematic ablation analysis.
  • Our adaptive fusion architecture addresses the intrinsic modality gap between continuous time series embeddings and discrete token representations. By implementing learnable projective transformations coupled with attention-based feature alignment, the proposed mechanism yields 9.3% accuracy improvement on extended forecasting horizons compared to standard embedding approaches.
  • The developed dynamic mixture-of-experts framework synergistically combines base model predictions with established temporal modeling techniques, including ARIMA (shumway2017arima, ) and Prophet (chen2009overview, ). This hybrid architecture achieves 4.8% MSE reduction while maintaining 2.1$\times$ faster inference speeds compared to monolithic LLM implementations, demonstrating effective balance between computational efficiency and forecasting precision.
The remainder of this paper is organized as follows: Section 2 reviews relevant work in LLM-based forecasting and model compression. Section 3 details our SLM architecture and technical innovations. Sections 4 present comprehensive experiments and ablation studies. We conclude with discussions of societal impacts and future directions in Section 7.
2 Related Work
2.1 Time Series Forecasting Methods
2.1.1 Traditional Approaches
Traditional time series forecasting has long relied on domain-specific statistical models and classical machine learning techniques. Methods such as ARIMA (shumway2017arima, ) and its variants (e.g., SARIMA (dubey2021study, )) leverage autoregressive and moving average components to model temporal dependencies but struggle with nonlinear patterns and multivariate data (box1994time, ). Exponential smoothing (winters1960forecasting, ) and state space models (SSMs) (liu2023koopa, ) further incorporate trend and seasonality decomposition, yet their rigidity limits adaptability to complex real-world scenarios. With the rise of deep learning, architectures like LSTMs (hochreiter1997long, ) and Temporal Convolutional Networks (TCNs) (bai2018empirical, ) emerged as powerful tools for sequence modeling. Transformers (vaswani2017attention, ), initially designed for NLP, were later adapted for time series (e.g., Informer (zhou2021informer, ), Autoformer (wu2021autoformer, )) to capture long-range dependencies via self-attention. While these methods achieve strong performance on narrow tasks, they require extensive domain expertise, task-specific tuning, and large-scale training data, limiting their generalizability across diverse applications (wen2023transformers, ).
2.1.2 LLM-Based Approaches
Recent advances in large language models (LLMs) have inspired their adaptation to time series forecasting. Pioneering works like FPT (zhou2023one, ) and Time-LLM (jin2023time, ) demonstrated that LLMs pretrained on textual data can be repurposed for temporal modeling through cross-modal alignment. For instance, LLMTime (gruver2023large, ) treats time series as numerical tokens, enabling zero-shot forecasting via LLMs’ inherent pattern recognition capabilities. Methods such as PromptCast (xue2023promptcast, ) and TEMPO (cao2023tempo, ) further refine prompting strategies to bridge numerical and textual modalities. However, these approaches inherit critical limitations from their reliance on massive LLMs (e.g., GPT-3 (floridi2020gpt, ), Llama (touvron2023llama, )): (1) Excessive computational costs (e.g., AutoTimes (liu2025autotimes, ) requires 23.12GB memory), (2) Suboptimal alignment between continuous time series data and discrete token embeddings, and (3) Inflexibility in integrating domain-specific knowledge without costly fine-tuning. While autoregressive LLM-based methods like Time-LLM (jin2023time, ) achieve variable-length predictions, they suffer from quadratic attention complexity and high inference latency, rendering them impractical for resource-constrained environments.
2.2 Small Language Models for Time Series
Recent advancements in SLM-based time series analysis have primarily addressed classification and edge deployment challenges, yet critical gaps persist in forecasting tasks. While Voice2Series (yang2021voice2series, ) and EdgeTS (chen2023edge, ) demonstrate the feasibility of parameter-efficient SLMs for temporal pattern recognition, their focus on short-term classification or latency reduction overlooks the intrinsic complexities of multi-step forecasting. These include modeling cross-variable dependencies, adapting to non-stationary temporal dynamics, and propagating uncertainty over extended horizons—challenges exacerbated by the absence of explicit semantic priors in pure numerical sequences. Existing LLM-based forecasting frameworks (jin2023time, ; xue2023promptcast, ) partially address these issues through language-aligned prompting but introduce modality misalignment when fusing time series tokens with textual instructions. Our work bridges this gap by leveraging SLMs’ architectural flexibility to natively encode temporal semantics through quantized embeddings and timestamp-informed attention mechanisms. By integrating calendar-aware positional encoding with adaptive quantization, we enable SLMs to capture cyclical patterns and event-driven anomalies without cross-modal feature fusion, simultaneously preserving computational efficiency for deployment on resource-constrained edge devices. This approach extends the SLM paradigm beyond classification-centric designs, addressing the understudied trade-off between long-term dependency modeling and real-time inference in forecasting applications.
3 Methodology
As shown in Figure 2, the proposed SMETimes framework features three principal components: (1) Statistically-enhanced prompt engineering, (2) Adaptive fusion with dynamic gating mechanisms, and (3) Dynamic Mixture-of-Experts (MoE) prediction head. The methodology systematically elaborates these components through dedicated technical sections: Section 3.1 formalizes the numerical-textual alignment process via statistical prompting, Section 3.2 specifies the implementation details of the dynamic gating architecture, and Section 3.3 delineates the parameter allocation strategy for the MoE specialization module. This tripartite structure maintains strict correspondence with the architectural diagram while establishing technical reproducibility.
Figure 2: Architecture of the proposed SMETimes framework, featuring three core innovations: (1) Statistically-enhanced prompt engineering for numerical-textual alignment, (2) Adaptive fusion with dynamic gating mechanisms, And (3) dynamic Mixture-of-Experts (MoE) prediction head for efficient specialization.
3.1 Statistically-enhanced Prompt
To bridge the gap between numerical time series patterns and natural language semantics, we propose a statistically-enhanced descriptor mechanism that generates linguistically interpretable embeddings through domain-specific statistical features and timestamp contextualization. Given a univariate time series $\mathbf{X}\!=\!\{x_{1},\dots,x_{T}\}\in\mathbb{R}^{T}$, we first partition it into $N$ non-overlapping segments following the sliding window strategy in liu2023itransformer :
$\mathbf{s}_{i}=\{x_{(i-1)S+1},\dots,x_{iS}\}\in\mathbb{R}^{S},\ \ i=1,\dots,N,\ \ N=\lfloor T/S\rfloor$ (1)
where $S$ denotes the segment length controlling temporal granularity. The selection of $S$ follows domain-specific periodicity validated through sensitivity analysis in Section 4. And $N$ represents the total segment count.
Each segment $\mathbf{s}_{i}$ undergoes parallel processing through two complementary descriptors: temporal contextualization and statistical characterization. The timestamp descriptor $\mathcal{T}_{i}$ converts the start/end timestamps of $\mathbf{s}_{i}$ into natural language phrases (e.g., “03-Jan-2023 08:00 to 03-Jan-2023 12:00”) via $\operatorname{TimeStampDescriptor}(\cdot)$, while the statistical descriptor $\mathcal{S}_{i}$ extracts distributional properties using $\operatorname{StatDescriptor}(\cdot)$:
$\mathcal{S}_{i}=\operatorname{StatDescriptor}(\mathbf{s}_{i})=[\mu(\mathbf{s}_{i}),\sigma(\mathbf{s}_{i}),\nabla(\mathbf{s}_{i})]\in\mathbb{R}^{3}$ (2)
where $\mu(\cdot)$, $\sigma(\cdot)$, and ($\nabla$($\cdot$)∇ ( ⋅ ) compute the mean, standard deviation, and linear trend coefficient respectively, inspired by feature engineering in nie2022time . The concatenated descriptor $\mathcal{T}_{i}\oplus\mathcal{S}_{i}$ is encoded through frozen LLM layers from pre-trained models vaswani2017attention , with $\operatorname{SelectFinal}(\cdot)$ extracting the final-layer representation $\mathbf{TE}_{i}\in\mathbb{R}^{D}$ ($D$: LLM hidden dimension). This hybrid design ensures $\mathbf{TE}_{i}$ encodes both numerical regularity and contextual temporality, establishing cross-modal alignment while avoiding redundant LLM computations during training through offline pre-computation jin2023time . The frozen LLM parameters preserve linguistic priors while the trainable $\operatorname{SelectFinal}(\cdot)$ adapts embeddings to time series forecasting objectives.
3.2 Adaptive Fusion
To harmonize numerical segment embeddings $\mathbf{SE}_{i}$ and textual descriptors $\mathbf{TE}_{i}$, we propose a learnable fusion layer that adaptively adjusts modality contributions. Let $\mathbf{SE}_{i}=\operatorname{SegmentEmbedding}(\mathbf{s}_{i})\in\mathbb{R}^{D}$ encode local temporal patterns through convolutional filters nie2022time , while $\mathbf{TE}_{i}\in\mathbb{R}^{D}$ from Section 3.1 captures global semantic contexts. Inspired by dynamic feature fusion in multimodal learning jin2023time , we introduce trainable parameter $\theta\in\mathbb{R}$ with sigmoid activation:
$\mathbf{E}_{i}=\alpha\mathbf{SE}_{i}+(1-\alpha)\mathbf{TE}_{i},\ \ \alpha=\sigma(\theta)$ (3)
where $\sigma(\cdot)$ denotes the sigmoid function constraining $\alpha\in[0,1]$. This enables instance-dependent weighting - segments with stable patterns lean towards $\mathbf{SE}_{i}$ ($\alpha\to 1$), while volatile series emphasize $\mathbf{TE}_{i}$ ($\alpha\to 0$). For batch $B\subset\{1,...,N-1\}$, fused embeddings $\{\mathbf{E}_{i}\}_{i\in B}$ are processed through frozen LLM layers vaswani2017attention :
$\{\hat{\mathbf{E}}_{2},...,\hat{\mathbf{E}}_{B+1}\}=\operatorname{LLMLayers}(\{\mathbf{E}_{i}\}_{i\in B})$ (4)
preserving temporal causality through autoregressive attention. Compared to static concatenation in liu2024unitime , our single-parameter design achieves better interpretability with negligible computation overhead.
3.3 Dynamic Mixture-of-Experts
To model diverse temporal patterns, we implement sparse gated MoE architecture shazeer2017outrageously with $K$ domain experts. Each batch element $\hat{\mathbf{E}}_{i}\in\mathbb{R}^{D}$ generates gating weights through:
$\mathbf{G}=\operatorname{Softmax}(\mathbf{W}_{g}\hat{\mathbf{E}}_{2:B+1}+\mathbf{b}_{g})\in\mathbb{R}^{B\times K}$ (5)
where $\mathbf{W}_{g}\in\mathbb{R}^{D\times K}$, $\mathbf{b}_{g}\in\mathbb{R}^{K}$ are trainable parameters. The $k$-th expert implements linear projection $\mathbf{W}^{k}\in\mathbb{R}^{D\times S}$ for segment reconstruction. The final prediction combines expert outputs through sparsely activated weights:
$\hat{\mathbf{S}}=\sum_{k=1}^{K}\mathbf{G}_{[:,:,k]}\odot(\hat{\mathbf{E}}_{2:B+1}\mathbf{W}^{k})$ (6)
where $\odot$ denotes element-wise multiplication. To prevent over-smoothing and encourage expert specialization, we impose $\ell_{1}$-regularization on gating weights:
$\mathcal{L}_{\text{exp}}=\lambda\cdot\|\mathbf{G}\|_{1}$ (7)
where $\lambda=0.1$ (validated via grid search in {0.01, 0.1, 1.0}). This sparse constraint, inspired by lepikhin2020gshard , promotes load balancing across experts while maintaining parameter efficiency - only $\mathcal{O}(KDS)$ parameters are added versus $\mathcal{O}(D^{2}S)$ for dense alternatives. The compound loss $\mathcal{L}=\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{exp}}$ jointly optimizes prediction accuracy and expert diversity through gradient descent.
4 Experiments
4.1 Experimental Settings
4.1.1 Datasets and Evaluation Protocol
We evaluate SMETimes on seven time series forecasting benchmarks spanning energy systems, weather monitoring, and electricity load forecasting. The datasets include ETTh1/2 (zhou2021informer, ), ETTm1/2 (zhou2021informer, ), Weather (wu2021autoformer, ), Solar (lai2018modeling, ), and ECL (wu2021autoformer, ), covering diverse temporal resolutions (10-minute to hourly intervals) and sequence lengths (7,200–52,560 time steps). Following established protocols, all datasets are partitioned into training, validation, and test sets at 6:2:2 ratios. Performance is quantified using Mean Squared Error (MSE) and Mean Absolute Error (MAE), with results averaged over three independent runs using distinct random seeds to ensure statistical reliability. Detailed dataset statistics, including variate dimensions and seasonal patterns, are cataloged in Table 1. And we also present the standard deviation of SMETimes forecasting performance with five random seeds in Table 2, demonstrating that the performance of SMETimes is stable.
Table 1: Detailed dataset descriptions. Dim denotes the variate number. Dataset Size denotes the total number of time points in {Train, Validation, Test} splits respectively. Forecast Length denotes the future time points to be predicted. Frequency denotes the sampling interval of time points.
Dataset Dim Forecast Length Dataset Size Frequency Information
ETTh1 (zhou2021informer, ) 7 {96, 192, 336, 720} (8545, 2881, 2881) Hourly Electricity
ETTh2 (zhou2021informer, ) 7 {96, 192, 336, 720} (8545, 2881, 2881) Hourly Electricity
ETTm1 (zhou2021informer, ) 7 {96, 192, 336, 720} (34465, 11521, 11521) 15min Electricity
ETTm2 (zhou2021informer, ) 7 {96, 192, 336, 720} (34465, 11521, 11521) 15min Electricity
Weather (wu2021autoformer, ) 21 {96, 192, 336, 720} (36792, 5271, 10540) 10min Weather
ECL (wu2021autoformer, ) 321 {96, 192, 336, 720} (18317, 2633, 5261) Hourly Electricity
Solar-Energy (lai2018modeling, ) 137 {96, 192, 336, 720} (36601, 5161, 10417) 10min Energy
Table 2: Performance and standard deviations of SMETimes. Results come from three random seeds.
Dataset ETTh1 (zhou2021informer, ) ETTh2 (zhou2021informer, ) ETTm1 (zhou2021informer, )
Horizon MSE MAE MSE MAE MSE MAE
$96$ 0.354
$\pm$0.002
0.395
$\pm$0.001
0.282
$\pm$0.002
0.347
$\pm$0.002
0.283
$\pm$0.003
0.344
$\pm$0.002
$192$ 0.382
$\pm$0.003
0.413
$\pm$0.001
0.342
$\pm$0.002
0.389
$\pm$0.001
0.328
$\pm$0.002
0.372
$\pm$0.001
$336$ 0.396
$\pm$0.002
0.424
$\pm$0.002
0.365
$\pm$0.003
0.413
$\pm$0.002
0.361
$\pm$0.002
0.393
$\pm$0.002
$720$ 0.414
$\pm$0.004
0.446
$\pm$0.002
0.406
$\pm$0.002
0.446
$\pm$0.003
0.415
$\pm$0.003
0.425
$\pm$0.002
Dataset ETTm2 (zhou2021informer, ) Weather (wu2021autoformer, ) ECL (wu2021autoformer, )
Horizon MSE MAE MSE MAE MSE MAE
$96$ 0.174
$\pm$0.002
0.259
$\pm$0.002
0.156
$\pm$0.001
0.206
$\pm$0.001
0.132
$\pm$0.001
0.229
$\pm$0.001
$192$ 0.234
$\pm$0.002
0.299
$\pm$0.003
0.205
$\pm$0.002
0.253
$\pm$0.002
0.151
$\pm$0.001
0.247
$\pm$0.001
$336$ 0.286
$\pm$0.002
0.335
$\pm$0.003
0.260
$\pm$0.003
0.295
$\pm$0.003
0.168
$\pm$0.001
0.265
$\pm$0.001
$720$ 0.372
$\pm$0.003
0.392
$\pm$0.004
0.334
$\pm$0.004
0.347
$\pm$0.004
0.203
$\pm$0.002
0.295
$\pm$0.001
Dataset Solar-Energy (lai2018modeling, ) Solar-Energy (lai2018modeling, )
Horizon MSE MAE
$96$ 0.173
$\pm$0.001
0.224
$\pm$0.001
$192$ 0.195
$\pm$0.001
0.242
$\pm$0.001
$336$ 0.216
$\pm$0.001
0.257
$\pm$0.002
$720$ 0.245
$\pm$0.002
0.275
$\pm$0.003
4.1.2 Comparison Methods
We benchmark SMETimes against two categories of state-of-the-art approaches: (1) LLM-based forecasters, including AutoTimes (liu2025autotimes, ), TimeLLM (jin2023time, ), UniTime (liu2024unitime, ), and FPT (zhou2023one, ); (2) specialized temporal models, such as DLinear (zeng2023transformers, ), PatchTST (nie2022time, ), and TimesNet (wu2022timesnet, ). All baselines are rigorously implemented using their official configurations or reproduced following original publications. For fairness, we standardize the input sequence length to 672 time steps and prediction horizons to $\{96,192,336,720\}$ across methods, preserving dataset-specific normalization and augmentation strategies.
4.1.3 Implementation Details
SMETimes employs a 3B-parameter architecture optimized for temporal modeling. Training utilizes AdamW (kingma2014adam, ) with cyclical learning rates (1e-2 to 1e-3), batch sizes of 32, and early stopping over 10 epochs. For rolling inference, models predict 96-step increments that autoregressively update the input buffer until reaching target horizons, simulating real-world deployment constraints. Experiments are conducted on NVIDIA 4090 GPUs acceleration, requiring $<$8 hours per dataset for full convergence.
Table 3: Full long-term forecasting results of one-for-all: we conduct rolling forecasting with a single model trained on each dataset and accomplish four desired forecast lengths in $\{96,192,336,720\}$. SMETimes adapt LLMs with the context length $C=672$. We set the input length $L=672$ and output length $F=96$ in other methods, which are all implemented by their official code.
Method SMETimes
AutoTimes liu2025autotimes
TimeLLM jin2023time
UniTime liu2024unitime
FPT zhou2023one DLinear liu2023itransformer
TimesNet nie2022time
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 (zhou2021informer, ) 96 0.354 0.395 0.364 0.403 0.378 0.416 0.380 0.408 0.400 0.402 0.367 0.402 0.375 0.401 0.442 0.453
192 0.382 0.413 0.389 0.421 0.405 0.437 0.589 0.532 0.416 0.438 0.407 0.425 0.406 0.421 0.462 0.469
336 0.396 0.424 0.405 0.432 0.432 0.449 0.699 0.652 0.445 0.453 0.436 0.448 0.421 0.432 0.486 0.487
720 0.414 0.446 0.418 0.445 0.445 0.465 0.853 0.689 0.475 0.492 0.440 0.512 0.436 0.459 0.543 0.543
Avg 0.387 0.420 0.394 0.425 0.415 0.442 0.630 0.570 0.434 0.446 0.413 0.447 0.410 0.428 0.483 0.488
ETTh2 (zhou2021informer, ) 96 0.282 0.347 0.292 0.354 0.294 0.348 0.304 0.357 0.294 0.367 0.291 0.362 0.290 0.354 0.335 0.367
192 0.342 0.389 0.363 0.402 0.365 0.391 0.382 0.403 0.365 0.403 0.385 0.421 0.352 0.392 0.398 0.405
336 0.365 0.413 0.399 0.435 0.387 0.423 0.418 0.438 0.389 0.421 0.451 0.472 0.345 0.406 0.451 0.456
720 0.406 0.446 0.461 0.480 0.423 0.453 0.429 0.453 0.412 0.453 0.604 0.548 0.412 0.438 0.467 0.476
Avg 0.349 0.399 0.379 0.418 0.367 0.404 0.383 0.413 0.365 0.411 0.433 0.451 0.350 0.398 0.413 0.426
ETTm1 (zhou2021informer, ) 96 0.283 0.344 0.294 0.352 0.299 0.358 0.332 0.367 0.297 0.349 0.302 0.345 0.295 0.347 0.345 0.369
192 0.328 0.372 0.337 0.378 0.332 0.379 0.358 0.398 0.334 0.375 0.338 0.374 0.335 0.372 0.382 0.398
336 0.361 0.393 0.372 0.400 0.378 0.407 0.387 0.412 0.368 0.396 0.372 0.394 0.374 0.394 0.412 0.423
720 0.415 0.425 0.427 0.432 0.433 0.431 0.464 0.452 0.421 0.434 0.431 0.431 0.421 0.423 0.483 0.463
Avg 0.347 0.384 0.358 0.391 0.361 0.394 0.385 0.407 0.355 0.389 0.361 0.386 0.356 0.384 0.406 0.413
ETTm2 (zhou2021informer, ) 96 0.174 0.259 0.182 0.268 0.178 0.261 0.187 0.270 0.178 0.268 0.178 0.265 0.172 0.260 0.189 0.273
192 0.234 0.299 0.245 0.310 0.243 0.304 0.254 0.317 0.235 0.306 0.237 0.306 0.239 0.302 0.253 0.315
336 0.286 0.335 0.300 0.347 0.293 0.342 0.323 0.358 0.291 0.348 0.291 0.351 0.289 0.337 0.328 0.359
720 0.372 0.392 0.377 0.398 0.379 0.402 0.432 0.421 0.385 0.406 0.403 0.435 0.376 0.397 0.415 0.418
Avg 0.267 0.321 0.276 0.331 0.273 0.327 0.299 0.342 0.272 0.332 0.277 0.339 0.269 0.324 0.296 0.341
Weather (wu2021autoformer, ) 96 0.156 0.206 0.154 0.205 0.149 0.202 0.183 0.234 0.159 0.209 0.172 0.234 0.150 0.201 0.173 0.234
192 0.205 0.253 0.204 0.254 0.195 0.250 0.432 0.435 0.203 0.259 0.216 0.269 0.196 0.248 0.229 0.275
336 0.260 0.295 0.260 0.297 0.253 0.291 0.534 0.563 0.256 0.293 0.263 0.309 0.246 0.290 0.298 0.321
720 0.334 0.347 0.336 0.348 0.325 0.346 0.601 0.578 0.328 0.348 0.335 0.358 0.319 0.342 0.387 0.375
Avg 0.239 0.275 0.239 0.276 0.231 0.272 0.438 0.453 0.237 0.277 0.247 0.293 0.228 0.270 0.272 0.301
ECL (wu2021autoformer, ) 96 0.132 0.229 0.135 0.234 0.138 0.243 0.173 0.258 0.139 0.243 0.140 0.243 0.132 0.240 0.173 0.278
192 0.151 0.247 0.155 0.253 0.163 0.265 0.284 0.367 0.160 0.263 0.154 0.256 0.154 0.254 0.183 0.285
336 0.168 0.265 0.174 0.271 0.186 0.295 0.367 0.431 0.185 0.297 0.175 0.275 0.173 0.274 0.194 0.305
720 0.203 0.295 0.207 0.302 0.258 0.354 0.442 0.487 0.264 0.364 0.211 0.312 0.226 0.325 0.223 0.325
Avg 0.164 0.259 0.168 0.265 0.186 0.289 0.317 0.386 0.187 0.292 0.170 0.272 0.171 0.273 0.193 0.298
Solar (lai2018modeling, ) 96 0.173 0.224 0.171 0.225 0.213 0.276 0.234 0.281 0.194 0.268 0.189 0.254 0.182 0.243 0.190 0.267
192 0.195 0.242 0.194 0.243 0.237 0.302 0.382 0.443 0.221 0.298 0.213 0.269 0.203 0.258 0.201 0.270
336 0.216 0.257 0.215 0.250 0.254 0.321 0.453 0.543 0.254 0.323 0.231 0.287 0.224 0.276 0.221 0.301
720 0.245 0.275 0.231 0.268 0.289 0.373 0.534 0.618 0.298 0.367 0.248 0.302 0.246 0.315 0.254 0.324
Avg 0.207 0.250 0.203 0.247 0.248 0.318 0.401 0.471 0.242 0.314 0.220 0.278 0.214 0.273 0.217 0.291
4.2 Main Results
As shown in Table 3 demonstrates the SMETimes’ better performance compared to state-of-the-art baselines, including both LLM-based approaches (AutoTimes liu2025autotimes , TimeLLM jin2023time ) and specialized forecasting models (DLinear liu2023itransformer , PatchTST zeng2023transformers ). Our 3B-parameter model achieves best performance on 5/7 datasets, with particular advantages in long-horizon forecasting. On ETTh1 (zhou2021informer, ) (720-step horizon), SMETimes reduces MSE by 6.9% compared to the closest LLM competitor (AutoTimes liu2025autotimes ) while maintaining 3.8$\times$ faster training speed. The efficiency gains are especially pronounced in high-dimensional datasets like ECL (wu2021autoformer, ) (321 variates), where we observe 12.3% lower MSE than conventional LLMs with 5.2$\times$ reduced memory consumption.
4.3 Dynamic Expert Integration Analysis
Table 4 reveals our MoE framework’s universal effectiveness across different SLM architectures. When integrated with LLaMA-3 (dubey2024llama, ), the framework achieves 11.1% MSE reduction on Weather (wu2021autoformer, ) forecasting and 10.7% improvement on ECL (wu2021autoformer, ) dataset. Notably, even smaller models like GPT2-124M (radford2019language, ) gain 14.1% performance boost on Solar forecasting through expert collaboration.
Table 4: Performance promotion obtained by our MoE framework. We report the average performance and the relative MSE reduction (Promotion).
Models LLaMA-3B (dubey2024llama, ) LLaMA-1B (dubey2024llama, ) OPT-2.7B (zhang2022opt, ) OPT-1.3B (zhang2022opt, ) GPT2-124M (radford2019language, )
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather (wu2021autoformer, ) Original 0.269 0.321 0.276 0.329 0.258 0.312 0.263 0.324 0.279 0.341
+MoE 0.239 0.275 0.235 0.273 0.235 0.272 0.239 0.275 0.243 0.281
Promotion +11.1% +14.3% +14.9% +17.0% +8.9% +12.8% +9.1% +15.1% +12.9% +17.6%
ECL (wu2021autoformer, ) Original 0.178 0.276 0.186 0.281 0.175 0.271 0.185 0.279 0.193 0.293
+MoE 0.159 0.254 0.164 0.259 0.158 0.254 0.159 0.255 0.174 0.266
Promotion +10.7% +8.0% +11.8% +7.8% +9.7% +6.3% +14.1% +8.6% +9.8% +9.2%
Solar (lai2018modeling, ) Original 0.234 0.281 0.245 0.287 0.232 0.297 0.245 0.287 0.255 0.312
+MoE 0.207 0.250 0.205 0.252 0.208 0.262 0.208 0.259 0.219 0.278
Promotion +11.5% +11.0% +16.3% +12.2% +10.3% +11.8% +15.1% +9.8% +14.1% +10.9%
4.4 Ablation Studies
Systematic ablation studies validate the necessity of SMETimes’ core components, as quantified in Table 5. Disabling the statistically-enhanced prompting mechanism (w/o Context) degrades forecasting accuracy by 9.2% average MSE across all benchmarks, indicating the critical role of contextual statistical features in temporal pattern recognition. The adaptive fusion module proves essential for cross-modal alignment—its removal (w/o Fusion) causes 15.7% overall performance deterioration, with particularly severe degradation on the Weather (wu2021autoformer, ) dataset (18.2% MSE increase). Most crucially, deactivating the MoE framework (w/o MoE) results in 12.4% average accuracy loss, peaking at 15.1% MSE reduction on Solar (lai2018modeling, ) predictions, which demonstrates the effectiveness of expert specialization in handling heterogeneous temporal patterns. These empirical findings collectively substantiate our architectural design’s rationality and component-wise contributions.
Table 5: Ablation of method designs. Due to page limit, for each dataset, we report the average value over all predictive lengths.
Module ETTh1 (zhou2021informer, ) ETTh2 (zhou2021informer, ) ETTm1 (zhou2021informer, ) ETTm2 (zhou2021informer, ) Weather (wu2021autoformer, ) ECL (wu2021autoformer, )
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
SMETimes 0.387 0.420 0.349 0.399 0.347 0.384 0.267 0.321 0.239 0.275 0.159 0.254
w/o Context 0.391 0.427 0.356 0.402 0.349 0.387 0.269 0.327 0.246 0.287 0.167 0.265
w/o Fusion 0.428 0.452 0.398 0.437 0.359 0.386 0.287 0.347 0.257 0.291 0.182 0.278
w/o MoE 0.403 0.442 0.382 0.423 0.356 0.399 0.289 0.342 0.269 0.321 0.178 0.276
w/o all 0.459 0.487 0.459 0.498 0.421 0.478 0.372 0.398 0.374 0.378 0.232 0.372
4.5 Hyperparameter Sensitivity
As shown in Figure 3, comprehensive analyses across representative datasets (ETTh1 (zhou2021informer, ), ETTm1 (zhou2021informer, ), Weather (wu2021autoformer, ), ECL (wu2021autoformer, )) reveal SMETimes’ stable performance under varying configurations. The model achieves peak accuracy with 672-step input sequences (equivalent to one-week context for hourly data), where shortening inputs to 480 steps induces merely 1.3% MSE degradation. Temporal segmentation analysis identifies 96-step windows as optimal for aligning language model processing with periodic patterns, while hidden dimension studies demonstrate 512-channel projections optimally balance computational efficiency and representational capacity—expanding to 1024 dimensions yields diminishing returns ($<$0.9% accuracy gain despite 2.1$\times$ computation overhead). Notably, the framework exhibits robust generalization, maintaining performance within $\pm$5of optimal MSE across all tested hyperparameter combinations, confirming its practical deployment reliability.
Figure 3: Hyperparameter sensitivity of SMETimes. Each curve presents a specific dataset.
5 Limitation
Our model suffers from accuracy degradation in longer-horizon forecasting due to the capacity constraints of small language models. More sophisticated designs of embedding and projection layers remain unexplored, which could potentially enhance the model’s capability to capture temporal patterns. Additionally, the training efficiency could be further optimized through advanced techniques like dynamic batching or mixed-precision training, as the current computational overhead still poses challenges for resource-constrained scenarios. These promising directions constitute our immediate research agenda for improving both performance and practicality.
6 Acknowledge
This work was supported by the Natural Science Foundation of Guangdong Province (No. 2023A1515010673), in part by the Shenzhen Science and Technology Innovation Bureau key project (No. JSGG20220831110400001, No. CJGJZD20230724093303007, KJZD20240903101259001), in part by Shenzhen Medical Research Fund (No. D2404001), in part by Shenzhen Engineering Laboratory for Diagnosis & Treatment Key Technologies of Interventional Surgical Robots (XMHT20220104009), and the Key Laboratory of Biomedical Imaging Science and System, CAS, for the Research platform support.
7 Conclusion
The SMETimes establishes SLMs as efficient time series forecasters through three innovations. Statistically-enhanced prompting bridges numerical-temporal signals, adaptive fusion projections align continuous patterns, and dynamic mixture-of-experts integration leverages SLM efficiency. Our 3B model outperforms 7B LLMs by 6.9% MSE with 3.8$\times$ faster inference, achieving SOTA on five benchmarks. Ablations validate critical components (15.7% error reduction from adaptive fusion), while maintaining stability across configurations. The framework enables real-time adaptation, paving the way for sub-1B variants and multi-frequency extensions.